# Characterizing the Post-Red Supergiant binary system AFGL 4106 and its complex nebula with SPHERE/VLT

### v1

@@ 3.4 Stars properties
With such luminosities and temperatures, those stars have radii of R1 = (209 12) R⊙ and R2 = (567 99) R⊙ at 3.19 kpc. This indicates that these stars are both supergiant in two different stages of evolution. The primary seems to be the more massive, the nebula was formed when it expelled its outer layers, leaving the star in the Post-RSG phase. The companion is likely in a less evolved stage, with a radius similar to giant stars, and may be surrounded with a possible expanding shell around the star caused by wind at the surface.
DUSTY computes both the inner radius of the dust shell and an estimate of the mass-loss rate. For our system, the inner radius is calculated as m, corresponding to 1.35 arcsec at a distance of 3.19 kpc. This radius is slightly less than the 1.6-1.9 arcsec derived by van Loon et al. (1999). To estimate the mass-loss rate, we applied the rescaling relation for dust-driven winds described by van Loon (2000)
Where is the gas-to-dust ratio assumed by Molster et al. (1999), and represents the bulk density of the dust. Using these parameters, we derive a total mass-loss rate of
@@ 4.1 Distance to AFGL 4106
The distance to AFGL 4106 is debated. Gaia’s parallax gives a distance of 0.94 0.26 kpc (Gaia Collaboration et al., 2016, 2023) but when applying the Bailer-Jones method (Bailer-Jones et al., 2021), using a probabilistic approach to estimate stellar distances and a prior constructed from a three-dimensional model of our Galaxy, we find a distance of 1.47 kpc. Due to its proximity with the Carina nebula Molster et al. (1999) assumed that AFGL 4106 was at a distance of 2.61 kpc
To improve the distance estimation, we combined the interstellar extinction measurement from Molster et al. (1999) with the value derived from GSP-Spec data (de Laverny, in preparation, private communication), and compared them with the extinction maps from G-TOMO (Lallement et al., 2022; Vergely et al., 2022). G-TOMO delivers three-dimensional extinction maps within 5 5 0.8 pc region centred on the Sun, using Gaia EDR3 and 2MASS photometry and parallaxes. This provides access to distance-dependent extinction curves in the solar neighbourhood, which can then be inverted: given the extinction, we can estimate a star’s distance. Taking an extinction value of mag, we derive a distance of 3.19 kpc.
As these are massive stars, their lifetime is particularly short - around 11 Myr for a 15 M⊙ star - so they are necessarily located close to the cluster where they were formed. The two closest star forming regions are the Carina nebula and the open cluster Westerlund 2, which are, respectively, at a projected distance of 2.1° and 1.8° of AFGL 4106 and at a distance of 2.61 kpc and 4.13 kpc. In the Carina’s reference frame, the proper motion is oriented in the direction of the South, with an absolute value of 4 mas yr-1 (Gaia Collaboration et al., 2023). However, the proper motion of AFGL 4106 is not consistent with an origin in the Carina Nebula, as shown in (Fig. 8). When re-referenced to the Carina reference frame, the proper motion vector does not point toward the nebula, indicating a different origin or kinematic history. In the Westerlund referential frame, the proper motion is also directed towards the south, with an absolute value of 4.5 mas yr-1. However, in this frame, the motion appears to be moving directly away from the cluster.
@@ 4.2 Radial profiles
The radial profiles of both stars were computed from ZIMPOL and IRDIS images, and the full width at half maximum (FWHM) was derived for each filter (Fig. 9). The measured FWHM values were then compared to the diffraction limit, (black line), and to the IRDIS and ZIMPOL modelled point spread function (PSF) generated using the Maoppy code (Fétick et al., 2019) (blue line).
@@ 4.3.1 HR diagram
Similarly, the secondary’s temperature and luminosity are consistent with those of a non-rotating star with an initial mass of about 9 M⊙, corresponding to an age of roughly 30 Myr and a current mass of 8.7 M⊙. Once again, this age is incompatible with that of Westerlund 2 and, more importantly, does not agree with the age inferred for the primary in the previous paragraph. This discrepancy further suggests that standard single-star evolutionary tracks may not adequately describe the system’s past.
@@
The secondary falls clearly within the RSG region, consistent with its classification. In contrast, the primary is located just below the so-called “Yellow Void.” It is too hot to be classified as an RSG and too faint to be considered a Yellow Hypergiant, placing it instead among the Yellow Supergiants (YSGs). Intermediate-mass stars can experience a ”blue loop” during their evolution, temporarily evolving from the RSG phase to warmer YSG or even Blue Supergiant (BSG) phases. However, the primary’s mass (M ¿ 12 M⊙) is likely too high for this mechanism to occur under standard evolutionary models. The system’s binarity is a factor that may alter its evolutionary pathway. However, the binary is too wide to have affected the evolution of the primary. For stars to truly affect each other they should be close, at most within a few au (Sana et al., 2012). This leaves open the possibility that AFGL 4106 experienced interaction with a third, close companion, or even a past merger, which might help explain a blue-loop-like excursion or its unusually low post-RSG luminosity.
@@ 4.3.2 Nebular shaping: influence of the companion
We have seen in Sec. 4.1 that the nebula shows signs of oscillation Fig. 5. These oscillations might come from the variation of the orbital speed of a companion, alongside the orbit curve (El Mellah et al., 2020). When the ratio between the wind speed and orbital velocity is low, the ejecta is constrained in the orbital plane. Thus, depending on the angle of observation, the shape of this ejecta might be elliptical, whereas when is larger, the ejecta is spherical, so that the nebula will appear spherical for all line of sights.
If the observed variation in ellipticity is attributed to the orbital motion of the binary, then, by applying Kepler’s third law with an oscillation period of approximately 1300-1950 years (depending on the assumed wind velocity of 20-30 km s-1, Sec 3.2.2), the resulting total system mass is estimated to be in the range of 175-400 M⊙. However, this estimate is inconsistent with the stellar parameters derived in the previous section. Moreover, such high total mass is improbable, as the formation of a binary system comprising two stars each exceeding 80 M⊙ is considered unlikely based on current star formation models.
@@
van Loon et al. (1999) reported radial-velocity variations with a period of 4500 days, interpreted as the orbital motion of the known binary and implying stellar masses M⊙, consistent with our results. If, instead, the radial velocity oscillation were attributed to an orbit involving a third companion located at 20 mas, the resulting total mass would be M⊙, which is unphysical. Matching the mass estimated in Sect. 4.3.1 would require a separation smaller than mas (i.e. au at 3.19 kpc).
@@
The shaping of the nebula is likely the result of a complex dynamic driven by, first, the influence of one or many companions, and second by a complex mechanism of ejection. It is also important to note that, the semi minor axis is almost aligned with the proper motion of the system. The interaction with the interstellar medium (ISM) may be responsible for the peculiar morphology of the nebula, particularly its large-scale elliptical shape with an eccentricity of approximately 0.5, as the stellar wind is decelerated by the ISM through the formation of a ”shock front”.
@@ 5 Conclusions
The main results of this study can be summarized as follows:
AFGL 4106 is binary system composed of a Post Red Supergiant and a Red Supergiant with a separation of 0.272 0.017 arcsec and temperature T1 = 6723 196 K and T2 = 3394 264 K.
The primary star appears slightly brighter than its companion, with a bolometric luminosity ratio of L1/L2 = 2.09, primarily attributed to its temperature being approximately 2 times higher.
Both stars have high bolometric luminosities with L1 = (7.9 0.18) L⊙ and L2 = (3.8 0.11) L⊙ with d in kpc.
The Radii R1 = (209 12) R⊙ and R2 = (567 99) R⊙ reveal that the stars are in two close but different stage of evolution, the primary in a post-RSG phase and the companion in a RSG phase.
Stellar evolution model gives an estimation of 15 M⊙ for the primary and 9 M⊙ for the secondary.
The proximity to the open cluster Westerlund 2, and the relative movement of both system, suggest that the star could be originated from the cluster. This give an upper bound on the age of the stars 2-3 Myr.
@@
The complex morphology of the nebula reveals a non isotropic and non continuous matter ejection from the system, as well as a strong interaction with the ISM.
This study of AFGL 4106 will enable us to constrain the dust and evolutionary models in this type of system, to better understand the formation and evolution of such nebulae. Further calculation and hydrodynamic model, combined with observation of the dynamics of the nebula, will help us to study the physics of such systems and the final step of star evolution.
@@ Acknowledgements.
This work is supported by the French National Research Agency (ANR) funded project PEPPER (ANR-20-CE31-0002).References

### v2

@@ 3.4 Star properties
With such luminosities and temperatures, those stars have radii of R1 = (209 12) R⊙ and R2 = (567 99) R⊙ at 3.19 kpc. This indicates that these stars are both supergiants in two different stages of evolution. The primary seems to be the more massive, and the nebula was formed when it expelled its outer layers, leaving the star in the post-RSG phase. The companion is likely in a less evolved stage, with a radius similar to giant stars, and may be surrounded by a possible expanding shell caused by wind at the surface.
DUSTY computes both the inner radius of the dust shell and an estimate of the mass-loss rate. For our system, the inner radius is calculated as m, corresponding to 1.35 arcsec at a distance of 3.19 kpc. This radius is slightly less than the 1.6-1.9 arcsec derived by van Loon et al. (1999). To estimate the mass-loss rate, we applied the rescaling relation for dust-driven winds described by van Loon (2000):
where is the gas-to-dust ratio assumed by Molster et al. (1999) and represents the bulk density of the dust. Using these parameters, we derive a total mass-loss rate of
@@ 4.1 Distance to AFGL 4106
The distance to AFGL 4106 is debated. Gaia’s parallax gives a distance of 0.94 0.26 kpc (Gaia Collaboration et al., 2016, 2023), but when applying the Bailer-Jones method (Bailer-Jones et al., 2021), using a probabilistic approach to estimate stellar distances and a prior constructed from a three-dimensional model of our Galaxy, we find a distance of 1.47 kpc. Due to its proximity with the Carina nebula Molster et al. (1999) assumed that AFGL 4106 was at a distance of 2.61 kpc
To improve the distance estimation, we combined the interstellar extinction measurement from Molster et al. (1999) with the value derived from GSP-Spec data (de Laverny, in preparation, private communication), and compared them with the extinction maps from G-TOMO (Lallement et al., 2022; Vergely et al., 2022). G-TOMO delivers three-dimensional extinction maps within a 5 5 0.8 pc region centred on the Sun, using Gaia EDR3 and 2MASS photometry and parallaxes. This provides access to distance-dependent extinction curves in the solar neighbourhood, which can then be inverted: given the extinction, we can estimate a star’s distance. Taking an extinction value of mag, we derive a distance of 3.19 kpc.
As these are massive stars, their lifetime is particularly short - around 11 Myr for a 15 M⊙ star - so they are necessarily located close to the cluster where they were formed. The two closest star forming regions are the Carina nebula and the open cluster Westerlund 2, which are, respectively, at a projected distance of 2.1° and 1.8° of AFGL 4106 and at a distance of 2.61 kpc and 4.13 kpc. In the Carina reference frame, the proper motion is oriented in the direction of the south, with an absolute value of 4 mas yr-1 (Gaia Collaboration et al., 2023). However, the proper motion of AFGL 4106 is not consistent with an origin in the Carina Nebula, as shown in (Fig. 8). When re-referenced to the Carina reference frame, the proper motion vector does not point towards the nebula, indicating a different origin or kinematic history. In the Westerlund referential frame, the proper motion is also directed towards the south, with an absolute value of 4.5 mas yr-1. However, in this frame, the motion appears to be moving directly away from the cluster.
@@ 4.2 Radial profiles
The radial profiles of both stars were computed from ZIMPOL and IRDIS images, and the full width at half maximum (FWHM) was derived for each filter (Fig. 9). The measured FWHM values were then compared to the diffraction limit, (black line), and to the IRDIS and ZIMPOL-modelled point spread function (PSF) generated using the Maoppy code (Fétick et al., 2019) (blue line).
@@ 4.3.1 HR diagram
Similarly, the secondary’s temperature and luminosity are consistent with those of a non-rotating star with an initial mass of about 9 M⊙, corresponding to an age of roughly 30 Myr and a current mass of 8.7 M⊙. Once again, this age is incompatible with that of Westerlund 2 and, more importantly, it does not agree with the age inferred for the primary in the previous paragraph. This discrepancy further suggests that standard single-star evolutionary tracks may not adequately describe the system’s past.
@@
The secondary falls clearly within the RSG region, consistent with its classification. In contrast, the primary is located just below the so-called yellow void. It is too hot to be classified as an RSG and too faint to be considered a yellow hypergiant, placing it instead among the yellow supergiants (YSGs). Intermediate-mass stars can experience a ’blue loop’ during their evolution, temporarily evolving from the RSG phase to warmer YSG or even blue-supergiant (BSG) phases. However, the primary’s mass (M ¿ 12 M⊙) is likely too high for this mechanism to occur under standard evolutionary models. The system’s binarity is a factor that may alter its evolutionary pathway. However, the binary is too wide to have affected the evolution of the primary. For stars to truly affect each other they should be close, at most within a few astronomical units (Sana et al., 2012). This leaves the possibility that AFGL 4106 experienced interaction with a third, close companion, or even a past merger, open, which might help explain a blue-loop-like excursion or its unusually low post-RSG luminosity.
@@ 4.3.2 Nebular shaping: Influence of the companion
We see in Sect. 4.1 that the nebula shows signs of oscillation (Fig. 5). These oscillations might come from the variation of the orbital speed of a companion, alongside the orbit curve (El Mellah et al., 2020). When the ratio between the wind speed and orbital velocity is low, the ejecta is constrained in the orbital plane. Thus, depending on the angle of observation, the shape of this ejecta might be elliptical, whereas when is larger, the ejecta is spherical, so the nebula will appear spherical for all lines of sight.
If the observed variation in ellipticity is attributed to the orbital motion of the binary, by applying Kepler’s third law with an oscillation period of approximately 1300-1950 years (depending on the assumed wind velocity of 20-30 km s-1, Sect. 3.2.2), the resulting total system mass is estimated to be in the range of 175-400 M⊙. However, this estimate is inconsistent with the stellar parameters derived in the previous section. Moreover, such high total mass is improbable, as the formation of a binary system comprising two stars each exceeding 80 M⊙ is considered unlikely based on current star formation models.
@@
van Loon et al. (1999) reported radial-velocity variations with a period of 4500 days, interpreted as the orbital motion of the known binary and implying stellar masses M⊙, which is consistent with our results. If, instead, the radial velocity oscillation were attributed to an orbit involving a third companion located at 20 mas, the resulting total mass would be M⊙, which is unphysical. Matching the mass estimated in Sect. 4.3.1 would require a separation smaller than mas (i.e. au at 3.19 kpc).
@@
The shaping of the nebula is likely the result of a complex dynamic driven by, first, the influence of one or many companions and, second, by a complex mechanism of ejection. It is also important to note that, the semi minor axis is almost aligned with the proper motion of the system. The interaction with the interstellar medium (ISM) may be responsible for the peculiar morphology of the nebula, particularly its large-scale elliptical shape with an eccentricity of approximately 0.5, as the stellar wind is decelerated by the ISM through the formation of a ’shock front’.
@@ 5 Conclusions
The main results of this study can be summarised as follows:
AFGL 4106 is a binary system composed of a post-RSG and an RSG with a separation of 0.272 0.017 arcsec and temperatures T1 = 6723 196 K and T2 = 3394 264 K.
The primary star appears slightly brighter than its companion, with a bolometric-luminosity ratio of L1/L2 = 2.09, primarily attributed to its temperature being approximately two times higher.
Both stars have high bolometric luminosities, with L1 = (7.9 0.18) L⊙ and L2 = (3.8 0.11) L⊙ with d in kpc.
The radii R1 = (209 12) R⊙ and R2 = (567 99) R⊙ reveal that the stars are in two close, but different, stages of evolution: the primary in a post-RSG phase and the companion in a RSG phase.
The stellar evolution model gives an estimation of 15 M⊙ for the primary and 9 M⊙ for the secondary.
The proximity to the open cluster Westerlund 2 and the relative movement of both systems suggest that the star could have originated from the cluster. This gives an upper bound on the age of the stars 2-3 Myr.
@@
The complex morphology of the nebula reveals a non-isotropic and non-continuous matter ejection from the system, as well as a strong interaction with the ISM.
This study of AFGL 4106 will enable us to constrain the dust and evolutionary models in this type of system, to better understand the formation and evolution of such nebulae. Further calculation and hydrodynamic models, combined with observations of the dynamics of the nebula, will help us to study the physics of such systems and the final step of star evolution.
@@ Acknowledgements.
This work was supported by the French National Research Agency (ANR) through the PEPPER project (ANR-20-CE31-0002). I would like to thank the referee for their constructive comments, which helped improve the quality of this paper.References